OpenStudy (anonymous):

x2-9=45 use square roots plz solve and show work!

OpenStudy (anonymous):

x squared -9=45 use square roots

OpenStudy (lilg132):

54 = x^2 square root 54 and you will have the answer for x

OpenStudy (anonymous):

what? can u explain how u got that!

OpenStudy (lilg132):

x^2 - 9 = 45 bring the numbers to one side x^2 = 45 + 9 (when you bring a number to the other side it changes to the opposite sign e.g. negative to positive) x^2 = 54 to get x by it self you need to square root it, but remember whatever you do to one side you also do to the other side so you square root 54 as well and you have the answer for x

OpenStudy (anonymous):

i get it thanks !

OpenStudy (lilg132):

what happened to clicking best answer :p

OpenStudy (anonymous):

hahah sry i was finishing anothwer problem on my sheet haha

OpenStudy (lilg132):

;)

OpenStudy (anonymous):

ANOTHER*